# DiffImp: Efficient Diffusion Model for Probabilistic Time Series Imputation with Bidirectional Mamba Backbone

### v1

@@ 3.1 Diffusion Models for Time Series Imputation
$p_{\theta}(X^{m}_{0:T}|X_{0},X_{o}^{c})=p(X_{T}^{m})\prod_{t=1}^{T}p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c}),$ (7)
@@
$p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c})=\mathcal{N}(X_{t-1}^{m};\mu_{\theta}(X_{t}^{m},t|X_{o}^{c}),\sigma_{\theta}(X_{t}^{m},t|X_{o}^{c})\bm{I}),$ (8)
the parameterized mean turns to:
$\mu_{\theta}(X_{t},t)=\frac{1}{\alpha_{t}}\left(X_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(X_{t},t|X_{o}^{c})\right),$ (9)
@@
$X_{t}=\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,$ (10)
and $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is a predefined variance scheduler and $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{t})$, hence we get the conditional diffusion loss for time series imputation task:
$\mathcal{L}=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(X_{t},t|X_{o}^{c})\|=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,t|X_{o}^{c})\|,$ (11)
@@
In the real world, the imputation problem encounters various complexities, such as different ratios of missing data, the positions of missing values within the sequence and the distribution of missing data. To simulate various complex missing situations in real-world scenarios, we adopt a self-supervised approach for training, i.e., applying a predefined mask to the complete dataset to construct corresponding dataset with missing data. We follow the same mask strategies in Alcaraz & Strodthoff (2023), including Random Missing (RM) which corresponds to the situation of uniformly random missing values, Random Block Missing (RBM) which corresponds to the situation of continuous missing values (missing intervals) in different channels and Blackout Missing (BM) which contains missing intervals at the same timestamps among different channels.
Figure 1: The self-supervised framework and training process of DiffImp. First, some of observed values are masked following the same missing pattern as the missing values (in red) to get masked targets ($X_{0}$, in magenta) and the condition input ($X_{o}^{c}$, in blue). The noisy input is obtain from $X_{0}$ and $\epsilon$ (in orange) sampled from $\mathcal{N}(0,I)$ The objective of the network is to minimize the difference between the parameterized noise $\epsilon_{\theta}(X_{t},t)$ and $\epsilon$. Solid lines in each time series represent observed values, while dashed lines represent missing values.
3.2 Model Architecture
The Overall Module Architecture Fig.1 illustrates the overall self-supervised framework and training process of our model. We first mask part of the observed values according to the pattern of missing values, where the masked values serve as the imputation target $X_{0}$ during training. The remaining observed values form the conditional input $X_{o}^{c}$ for the noise prediction network $\epsilon_{\theta}$. We then combine $X_{0}$ with noise $\epsilon$ sampled from a standard normal distribution to obtain the noisy input $X_{t}$. Both $X_{o}^{c}$, $X_{t}$, and the diffusion step $t$ are fed into the noise prediction network $\epsilon_{\theta}$ to get the parameterized noise. The network minimizes the difference between $\epsilon_{\theta}$ and $\epsilon$ according to Eq.11.
As shown in Fig.2, the forward process of $\epsilon_{\theta}$ are as follows: For each diffusion step, the input consists of the following parts: noisy input $X_{t}$, the condition input $X_{o}^{c}$ and the diffusion step $t$. To begin with, the inputs are embedded to the latent diffusion space. The embedding module of noisy inputs and condition inputs share a similar model structure, which consists of a linear projection module followed by an SMM block in Fig.3b. The SMM block is composed of stacks of Bidirectional Attention Mamba (BAM) blocks and Channel Mamba Blocks (CMB), which is introduced in the next part. Due to the relatively limited information from $t$, the embedding module of $t$ only consists of linear projection modules. After the embedding step, the embedded diffusion step is concatenated with the input embeddings. The concatenated embeddings are fed in to a SMM module. Then the output of the SMM module is concatenated with the condition embeddings. After feeding the final embeddings to another SMM module and final projection module, we can get the noise predictions $\epsilon_{\theta}(X_{t},t)$. The training and sampling algorithm is detailed in Alg.1 and Alg.2.
Figure 2: Architecture of $\epsilon_{\theta}$ in DiffImp
Mamba Encoders for Bidirectional Modeling For probabilistic time series imputation tasks, the objective is to attain a more precise posterior estimation for the missing points contingent upon the observed points. Therefore, our proposed module should achieve two key objectives: Firstly, it should possess bidirectional analysis capability, which means that the model should be able to capture dependencies in both the forward and reverse temporal directions. Secondly, considering that the known points at different positions relative to the missing point have varying distances, the model should assign different weights to difference timestamps. To address these issues, we devise a bidirectional attention Mamba module (BAM). BAM takes the representations from previous layers as input, which are then fed into two distinct PNM modules (Fig.3a), enabling the model to capture bidirectional dependencies. More specifically, temporal attention is implemented by assigning different values to various time steps in the sequence, where the temporal attention module also receives the previous layer’s representation and learns the weights for different timestamps. The details of BAM are shown in Fig.3d.
Mamba Encoders for Inter-channel Modeling In the context of multivariate time series, inter-dependencies exist among variables across different channels. The effective modeling of these inter-channel correlations is instrumental in capturing the intrinsic characteristics of the time series more adeptly. Additionally, when analyzing the relationships between channels, the order of the channels does not exhibit the sequential dependencies as that among timestamps. Consequently, we employ unidirectional channel dependency modelling architecture, termed as Channel Mamba Block (CMB). We first transpose the input time series representation for processing on the channel dimension. The transposed representations are then subjected to a normalization module and processed through a PNM block, yielding a more profound feature representation. The details of CMB are presented in Fig.3c.
Figure 3: Details of PNM, SMM, CMB, BAM block in the noise prediction module. (a) PNM: backbone module based on Mamba. (b) SMM: core components of noise prediction module, composed of stacks of BAM and CMB. (c) CMB: unidirectional module for inter-channel dependency modeling. (d) BAM: bidirectional module with temporal attention for intra-channel, multi-range dependency modeling.
Algorithm 1 Training Procedure of DiffImp
@@ 3.2 Model Architecture
1:  Input: Observed sequence $x_{0}$, number of iterations $N$, variance scheduler $\beta_{t}$
@@
Algorithm 2 Sampling Procedure of DiffImp
@@
Complexity Analysis While dealing with the input sequences, the core component of our module is the PNM module in Fig.3 and the self-attention module in the Transformer architecture, respectively. In this part, we will give a brief analysis about the time and space complexity in the SSM module and self-attention module [^1] We do not take the time and space complexity of MLPs before the self-attention module or SSM module into consideration.. The time complexity of self-attention module is $O(CL^{2})$ and the space complexity is $O(L^{2}+CL)$, where $L$ is the length of the input sequence and $C$ is the channel of the input sequence.
@@
In our method, the forward process described in Eq.2 is implemented by converting the process to multiplications of structured matrices, which is of time complexity $O(NCL)$ and of space complexity $O(CL+N(C+L))$ ($N$ is a constant number and set as $16$ by default). This indicates that our model is of linear time and space complexity with respect to the sequence length $L$, which ensures scalability and reduces memory cost for longer sequences.
4 Experiments
4.1 Experiment Settings
Datasets and Experimental Settings We conduct experiments on three real-world datasets to validate the effectiveness of our approach. These datasets span multiple domains, namely Electricity dataset Asuncion & Newman (2007), MuJoCo dataset Rubanova et al. (2019b) and ETTm1 dataset Zhou et al. (2021).
@@ 4.1 Experiment Settings
All experiments are conducted using PyTorch Paszke et al. (2019) in Python 3.9 and execute on an NVIDIA RTX3090 GPU. The training process is guided by Eq.11, employing the ADAM optimizer Kingma & Ba (2015) with a learning rate of $2\times 10^{-4}$. More details about the datasets and experimental settings can be found in the Appendix.
Evaluation Metrics and Baselines To achieve an extensive evaluation of imputation performance, diverse metrics are utilized for evaluating deterministic imputation results, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), and Root Mean Square Error (RMSE). Due to reproducibility reasons of baselines, we compare with different baselines and report different metrics for different datasets. The datasets and corresponding baseline and metrics are listed in Table.2. We follow the same settings and dataset preprocessings as Alcaraz & Strodthoff (2023) and collect all the baseline results from the same paper.
As for the evaluation of probabilistic imputation, we calculate the Continuous Ranked Probabilistic Score-sum (CRPS-sum) on the electricity dataset. The CRPS-sum results are collected from Yan et al. (2024). In all the tables of our experiment results, the best results are in bold and second best results are underlined. All the deterministic metrics are maintained by running the experiment for 3 times and CRPS-sum is obtained by 10 runs.
Table 2: Datasets and corresponding evaluation metrics and baselines for time series imputation and forecasting task.
Dataset Task Metric Baseline
Electricity Imputation MAE; RMSE; MRE
M-RNN Yoon et al. (2019); GP-VAE Fortuin et al. (2020);
BRITS Cao et al. (2018); SAITS Du et al. (2023); CSDI; SSSD
MuJoCo Imputation MSE
RNN GRU-D Che et al. (2018); ODE-RNN Rubanova et al. (2019a);
NeuralCDE Morrill et al. (2021); Latent-ODE Rubanova et al. (2019a);
NAOMI Liu et al. (2019); NRTSI Shan et al. (2023a); CSDI; SSSD
ETTm1 Forecasting MAE; MSE
LSTNet Lai et al. (2018); LSTM Bahdanau et al. (2015);
Reformer Kitaev et al. (2020); LogTrans Li et al. (2019);
Informer Zhou et al. (2021); Autoformer Wu et al. (2021); CSDI; SSSD
4.2 Time Series Imputation
@@
CSDI 2.4e-4$\pm$3e-5 6.1e-4$\pm$1.0e-4 4.84e-3$\pm$2e-5
SSSD 5.9e-4$\pm$8e-5 1e-3$\pm$5e-5 1.90e-3$\pm$3e-5
DiffImp (Ours) 2.7e-4$\pm$1e-5 3.16e-4$\pm$9.77e-6 6.5e-4$\pm$1e-4
Table.4 shows the experimental results on the Electricity dataset, where we apply the RM missing pattern with missing ratios of 10%, 30%, and 50%. We achieve the best results across all metrics with a 30% missing ratio, significantly outperforming other methods. In the experiments with 10% and 50% missing ratios, we obtain results with only a slight gap to the SOTA models.
Table 4: MAE and RMSE results on Electricity Dataset
@@
Model MAE RMSE MRE MAE RMSE MRE MAE RMSE MRE
M-RNN 1.244 1.867 66.6% 1.258 1.876 67.3% 1.283 1.902 68.7%
GP-VAE 1.094 1.565 58.6% 1.057 1.571 56.6% 1.097 1.572 58.8%
BRITS 0.847 1.322 45.3% 0.943 1.435 50.4% 1.037 1.538 55.5%
SAITS 0.735 1.162 39.4% 0.790 1.223 42.3% 0.876 1.377 46.9%
CSDI 1.510$\pm$3e-3 15.012$\pm$4e-2 81.10$\pm$1e-10.921$\pm$8e-3 8.732$\pm$7e-2 49.27$\pm$4e-10.278$\pm$4e-3 2.371$\pm$3e-2 14.93$\pm$1e-1%
SSSD 0.345$\pm$1e-4 0.554$\pm$5e-5 18.4$\pm$5e-30.407$\pm$5e-4 0.625$\pm$1e-4 21.8$\pm$00.532$\pm$1e-4 0.821$\pm$1e-4 28.5$\pm$1e-2%
DiffImp (Ours) 0.378$\pm$6e-4 0.522$\pm$3e-3 20.2$\pm$1e-20.348$\pm$1e-3 0.496$\pm$2e-3 18.6$\pm$1e-10.546$\pm$3e-3 0.837$\pm$7e-3 29.2$\pm$2e-1%
Probabilistic Imputation Results Table.5 presents a comparison of our method with other probabilistic time series imputation methods based on the CRPS-sum metric. The baselines for CRPS-sum include Tashiro et al. (2021); Chen et al. (2023); Alcaraz & Strodthoff (2023); Yan et al. (2024); Kollovieh et al. (2023). The experimental results show that our method achieves a 21.4% performance improvement compared to the second-best method. This indicates that our method models the data distribution of the sequence more accurately than other baseline methods.
Table 5: The CRPS-sum results on electricity dataset
Model CSDI CSBI SSSD TS-Diff D3M DiffImp(Ours)
CRPS-sum 2.14e-2$\pm$8e-3 2.19e-2$\pm$7e-3 1.96e-2$\pm$1e-3 2.23e-2$\pm$6e-3 1.92e-2$\pm$4e-3 1.51e-2$\pm$4e-4
4.3 Time Series Forecasting
As mentioned in 3.1, the probabilistic time series forecasting problem can be treated as a variant of the probabilistic time series imputation problem (as a special case of the missing manner BM). Therefore, we also conduct experiments to validate the effectiveness of our experiments on probabilistic time series forecasting task. Following the setup in previous works, we test five different forecasting horizons: 24, 48, 96, 288, and 672 time steps, with corresponding conditional lengths (i.e., the length of observed sequence) of 96, 48, 284, 288, and 384 time steps.
Table.6 presents the experimental results on the ETTm1 dataset. Our method achieves state-of-the-art performance on prediction length of 24 and 96, outperforms other imputation-based algorithms at the prediction length of 672, and shows only a slight gap compared to the best imputation-based algorithms at the prediction length of 48 and 288.
Table 6: MSE and MAE results on ETTm1 dataset
Forecasting Length 24 48 96 288 672
Model MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
LSTNet 1.170 1.968 1.215 1.999 1.542 2.762 2.076 1.257 2.941 1.917
LSTMa 0.629 0.621 0.939 1.392 0.913 1.339 1.124 1.740 1.555 2.736
Reformer 0.607 0.724 0.777 1.098 0.945 1.433 1.094 1.820 1.232 2.187
LogTrans 0.412 0.419 0.583 0.507 0.792 0.768 1.320 1.462 1.461 1.669
Informer 0.369 0.323 0.503 0.494 0.614 0.678 0.786 1.056 0.926 1.192
CSDI 0.370$\pm$3e-3 0.354$\pm$1.5e-2 0.546$\pm$2e-3 0.750$\pm$4e-3 0.756$\pm$1.1e-2 1.468$\pm$4.7e-2 0.530$\pm$4e-3 0.608$\pm$3.5e-2 0.891$\pm$3.7e-2 0.946$\pm$5.1e-2
Autoformer 0.403 0.383 0.453 0.454 0.463 0.481 0.528 0.634 0.542 0.606
SSSD 0.361$\pm$6e-3 0.351$\pm$9e-3 0.479$\pm$8e-3 0.612$\pm$2e-3 0.547$\pm$1.2e-2 0.538$\pm$1.3e-2 0.648$\pm$1.0e-2 0.797$\pm$5e-3 0.783$\pm$6.6e-2 0.804$\pm$4.5e-2
DiffImp (Ours) 0.282$\pm$1.8e-2 0.331$\pm$9.9e-3 0.679$\pm$5.6e-3 0.548$\pm$5.6e-4 0.3906$\pm$1.3e-2 0.4211$\pm$8.5e-3 0.621$\pm$2.1e-3 0.741$\pm$3.3e-3 0.683$\pm$3.1e-3 0.783$\pm$6.8e-3
4.4 Visualization Results
Figure 4: Visualized results of probabilistic time series imputation on MuJoCo dataset.
Fig.4 shows the visualization results for channel 5 and channel 7 on the MuJoCo dataset with a 90% missing ratio. From the figure, we can see that almost all ground truth values for the points to be imputed fall within the 95% confidence interval, and most of the ground truth values are within the 50% confidence interval, which demonstrates the effectiveness of our method.
4.5 Sampling Time Analysis
Table presents a comparison of sampling time between our method and other backbone-based methods across different datasets. We find that, in terms of sampling time, our method performs better compared to transformer-based method and previous SSM-based method, demonstrating the efficiency of our backbone choice. The data in the table represent the time of a single sample process.
Table 7: Sampling time of CSDI, SSSD and DiffImp on Electricity and MuJoCo dataset, - denotes the corresponding code is note provided in original papers.
CSDI SSSD DiffImp
Electricity 1.43s - 1.012s
MuJoCo - 4.16s 1.01s
4.6 Ablation Studies
To validate the effectiveness of the proposed module, we conduct ablation experiments on the following aspects: 1) the bidirectional modeling 2) the temporal attention mechanism 3) the inter-channel multivariate dependencies. All experiments are conducted on the MuJoCo dataset with the missing ratio 90%. During ablation experiments, we find out that our model converges much slower than other models in the ablation experiment, so we train till all models are converged (for same number of iterations, even if it has already been converged). The hyperparameters in the ablation studies are presented in the appendix.
The results are shown in Table.8. It can be observed that the module equipped with BAM and CMB block performs the best, significantly outperforming the results of removing any one of these components across all four metrics. The temporal attention module has the largest impact on the model, and its removal leads to a significant performance drop. Similarly, removing the CMB module also results in a notable degradation in performance. On the other hand, adjusting the BAM module to its unidirectional form also causes some degree of performance decrease. This fully demonstrates the effectiveness of our proposed blocks.
Table 8: Experimental results of Ablation Study

### v2

@@ 3.4. Diffusion Models for Time Series Imputation
(7) $p_{\theta}(X^{m}_{0:T}|X_{0},X_{o}^{c})=p(X_{T}^{m})\prod_{t=1}^{T}p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c}),$
@@
(8) $p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c})=\mathcal{N}(X_{t-1}^{m};\mu_{\theta}(X_{t}^{m},t|X_{o}^{c}),\sigma_{\theta}(X_{t}^{m},t|X_{o}^{c})\bm{I}),$
Then, the parameterized mean turns to:
(9) $\mu_{\theta}(X_{t},t)=\frac{1}{\alpha_{t}}\left(X_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(X_{t},t|X_{o}^{c})\right),$
@@
(10) $X_{t}=\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,$
$\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is a predefined variance scheduler, and $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{t})$.
Finally, we get the conditional diffusion loss for the time series imputation task:
(11) $\mathcal{L}=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(X_{t},t|X_{o}^{c})\|=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,t|X_{o}^{c})\|,$
@@
The network is trained based on Eq.11 to obtain the parameterized noise estimation function $\epsilon_{\theta}$. And the missing values are estimated by iterative sampling using $\epsilon_{\theta}$. The training and sampling algorithm is detailed in Alg.1 and Alg.2.
Compared with the attention-based backbones, the advantage of our proposed backbone lies in the following:
  1. (1)
[…]
  2. (2)
    Our proposed backbone has a controllable frequency response. From a frequency-domain perspective, SSMs offer controllable frequency responses $\hat{K}(\omega)=C(i\omega I-A)^{-1}B$, allowing them to suppress broadband noise while preserving signal components concentrated in specific frequency bands. This frequency selectivity enables more robust and adaptive representation learning, whereas attention mechanisms, lacking frequency bias, may fail under similar conditions.
Algorithm 1 Training Procedure of SSD-TS
@@
1:Input: Observed sequence $x_{0}$, number of iterations NN, variance scheduler $\beta_{t}$
@@
Algorithm 2 Sampling Procedure of SSD-TS
@@ 3.5. Complexity Analysis
In this part, we will give a brief analysis about the time and space complexity in the SSM module and self-attention module [^1] We do not take the time and space complexity of MLPs before the self-attention module or SSM module into consideration.. While dealing with the input sequences, the core component of our module is the PNM module in Fig.3 and the self-attention module in the Transformer architecture, respectively. The time complexity of self-attention module is $O(CL^{2})$ and the space complexity is $O(L^{2}+CL)$, where $L$ is the length of the input sequence and $C$ is the channel of the input sequence.
@@
In our method, the forward process described in Eq.2 is implemented by converting the process to multiplications of structured matrices, which is of time complexity $O(NCL)$ and of space complexity $O(CL+N(C+L))$ (NN is a constant number and set as $16$ by default). This indicates that our model is of linear time and space complexity with respect to the sequence length $L$, which ensures scalability and reduces memory cost for longer sequences.
4. Experiments
4.1. Experimental Settings and Evaluation Metrics
@@
All experiments are conducted using PyTorch (Paszke et al., 2019) in Python 3.9 and executed on an NVIDIA RTX3090 GPU. The training process is guided by Eq.11, employing the ADAM optimizer (Kingma and Ba, 2015) with a learning rate of $2\times 10^{-4}$. The hyperparameters in our experiments can be found in Appendix.7.1.2.
For time series modeling tasks, we explicitly categorize the problem into two paradigms based on the model’s capacity to quantify uncertainty, i.e., probabilistic time series modeling and deterministic time series modeling. To evaluate the performance of deterministic modeling methods, we report the Mean Absolute Error (MAE), the Mean Squared Error (MSE), Mean Relative Error (MRE) and Root Mean Squared Error (RMSE) between outputs of the model and ground-truth time series data. As defined in Definition.2.1, the original time series is denoted as $y\in\mathbb{R}^{K\times L}$, the output time series is denoted as $\hat{y}\in\mathbb{R}^{K\times L}$, $M$ is the indicator matrix, the metrics are formulated as follows:
Mean Absolute Error (MAE): MAE calculates the average $L_{1}$ distance between ground truth and the output values alongside the channel dimension, which is formulated as:
(12) $\mathbf{MAE}(y,\hat{y})=\frac{1}{k}\sum_{i=1}^{K}\sum_{j=1}^{L}|(y-\hat{y})\odot(1-M)|_{i,j}$
Mean Square Error (MSE): MSE calculates the average $L_{2}$ between ground truth and the output values alongside the channel dimension, which is formulated as:
(13) $\mathbf{MSE}(y,\hat{y})=\frac{1}{k}\sum_{i=1}^{K}\sum_{j=1}^{L}((y-\hat{y})\odot(1-M))^{2}_{i,j}$
Root Mean Square Error (RMSE): RMSE is the square root of the average $L_{2}$ between ground truth and the output values alongside the channel dimension:
@@
Mean Relative Error (MRE): MRE estimates the relative difference between $y$ and $\hat{y}$:
(15) $\mathbf{MRE}(y,\hat{y})=\frac{1}{k}\sum_{i=1}^{K}\sum_{j=1}^{L}(1-M)_{i,j}\odot\frac{|(y-\hat{y})|_{i,j}}{y_{i,j}}$
For probabilistic time series modeling evaluation, we additionally evaluate the Continuous Ranked Probability Score (CRPS). CRPS measures the integral squared difference between the predicted and empirical cumulative distribution functions (CDFs), providing a pointwise evaluation of probabilistic calibration. By applying CRPS, we can effectively evaluate the consistency between the estimated distribution and the real distribution. The CRPS metric is defined as follows:
Continuous Ranked Probabilistic Score (CRPS): Given an estimated probability distribution function $F$ modeled with an observation $x$, CRPS evaluates the compatibility and is defined as the integral of the quantile loss for all quantile levels:
(16) $\mathbf{CRPS}(F^{-1},x)=\int_{0}^{1}\Lambda_{\alpha}(F^{-1}(\alpha,x))\ \text{d}\alpha,$
where $\Lambda_{\alpha}(q,y)=(\alpha-\bm{1}_{y<q})(y-q),\alpha\in[0,1]$ and $\bm{1}_{y<q}$ the indicator function, i.e., if $y<q$, the value of the indicator function is $1$, else 0.
Following (Tashiro et al., 2021; Yan et al., 2024), we separate the interval $[0,1]$ to 2020 quantile levels with a stepsize of $s=0.05$, and the estimated value of CRPS is:
(17) $\mathbf{CRPS}(F^{-1},x)\approx\sum_{i=1}^{19}\frac{2\Lambda_{i\cdot s}(F^{-1}(i\cdot s,x))}{19}$
For the whole time series $X\in\mathbb{R}^{K\times L}$, the CRPS value is normalized for all time steps and channels:
(18) $\mathbf{CRPS}(F^{-1},X)=\frac{\sum_{i=1}^{K}\sum_{j=1}^{L}\mathbf{CRPS}(F^{-1}_{i,j},X_{i,j})}{\sum_{i=1}^{K}\sum_{j=1}^{L}|X_{i,j}|}$
4.2. Time Series Imputation
4.2.1. Datasets
To evaluate the performance of our model on time series imputation tasks, we employ three benchmark datasets, including the MuJoCo dataset (Rubanova et al., 2019), the Air Quality dataset (AQI) (Yi et al., 2016) and Physionet Challenge 2012 dataset (Silva et al., 2012). Please refer to Appendix.7.1.1 for more details about the datasets.
It is worth noting that we select these three datasets to include as many potential missing data scenarios as possible. For the Physionet dataset, consistent with previous studies (Cao et al., 2018; Suo et al., 2020; Qiu et al., 2024, 2025a), we process the dataset into hourly time series, each containing 48 time steps, with approximately 80% of the data missing. Since the dataset does not include ground truth, we randomly select 10%, 50%, and 90% of the data as ground truth in the test set. For the Air Quality dataset, around 13% of the data is missing in a non-random manner. As this dataset contains structured artificial ground truth, we do not test different missing ratios on the AQI dataset. The MuJoCo dataset, on the other hand, is a complete dataset and we apply hand-crafted masks to simulate masking patterns that are not present in the Physionet and Air Quality datasets.
4.2.2. Baselines
Correspondingly, in our experiments, the baselines include both deterministic and probabilistic imputation models. The deterministic models involve: BRITS (Cao et al., 2018), RDIS (Choi et al., 2023), SSGAN (Miao et al., 2021), TIDER (Liu et al., 2023), and SAITS (Du et al., 2023). The probabilistic imputation models encompass various distribution estimation methods, such as: methods based on variational autoencoders (V-RIN (Mulyadi et al., 2022)), methods based on diffusion models (CSDI (Tashiro et al., 2021), SSSD (Alcaraz and Strodthoff, 2023), D3M(Yan et al., 2024), TS-Diff (Liu et al., 2023)), methods based on Schrödinger Bridge (CSBI (Chen et al., 2023)), and methods based on Gaussian process (GP-VAE (Fortuin et al., 2020)). And on the MuJoCo dataset, we compare our models against RNN GRU-D (Che et al., 2018), ODE-RNN (Rubanova et al., 2019), NeuralCDE (Morrill et al., 2021), Latent-ODE (Rubanova et al., 2019), NAOMI (Marisca et al., 2022), CSDI (Tashiro et al., 2021) and SSSD (Alcaraz and Strodthoff, 2023) to be consistent with SSSD (Alcaraz and Strodthoff, 2023).
4.2.3. Experimental Results and Analysis
Table 2 presents the experimental results on the MuJoCo dataset for the random missing (RM) case, with different missing rate scenarios. Under the two higher missing rates (80% and 90%), our method achieves the best performance, with improvements of 48.2% (at 80% missing rate) and 65.8% (at 90% missing rate) compared to the second-best results. At 70% missing rate, our method ranks second, with only an 11.1% difference from the best-performing CSDI. This demonstrates the strong modeling capability of our method on datasets with high missing rates under random missing scenario.
@@ 4.2.3. Experimental Results and Analysis
[…]
Table.3 presents the RMSE performance of our method on two real-world datasets with missing data. We observe that on the Physionet dataset, our method achieves the best results across three different missing rates (10%, 50%, and 90%), with improvements of 22.6% (at 10% missing rate), 17.2% (at 50% missing rate), and 23.5% (at 90% missing rate). Additionally, on the AQI dataset, our method ranks second. These results demonstrate that our method outperforms other approaches in modeling time series data on real-world datasets with missing values.
Table 3. RMSE results on Physionet and Air quality dataset. The mean and standard error are obtained by 3 runs and the best result are in bold while the second best results are underlined.
Physionet AQI
10% missing 50% missing 90% missing
[…]
Table.4 presents the CRPS performance of our method on two real-world datasets with missing data. Compared to other probabilistic time series methods, our approach achieves the lowest CRPS scores in three out of the four tasks (10% and 50% on PhysioNet, and AQI). This indicates that, relative to other probabilistic time series modeling methods, our approach demonstrates the most accurate modeling capability of the distribution of unknown points within the sequence.
Table 4. CRPS results on Physionet and Air quality dataset. The mean and standard error are obtained by 3 runs and the best result are in bold while the second best results are underlined.
Physionet AQI
10% missing 50% missing 90% missing
[…]
In addition, we also compare the performance of diffusion models with different backbones. Existing imputation methods mainly cover two-kinds of backbones: transformer-based (Vaswani et al., 2017) backbones and state-space model-based (Gu et al., 2022) backbones. The core architecture of CSDI is a spatial-temporal transformer, which first embeds the input tokens and captures the inter- and intra- channel dependency with two transformer encoders. D3M adopts a similar attention mechanism as the backbone, with the addition of gating mechanisms and an exponential moving average (EMA) module. SSSD modifies the backbone of CSDI by replacing the spatial-temporal transformer with the S4 structure proposed in (Gu et al., 2022), which is also utilized in TS-Diff.
Table 5. Experimental Results of Ablation Study